# In-depth characterization of the Kepler-10 three-planet system with HARPS-N RVs and Kepler TTVs

### v1

@@ 3.5 Simultaneous modeling of radial velocities and transit timing variations
The TRADES code cannot manage complex stellar activity, it can only fit an RV offset ($\gamma$) and a jitter term ($\log_{2}{\sigma_{\rm j}}$) for each RV data-set; this is not an issue in our case, given the low level of stellar activity and the fact that the planet parameters do not critically depend on the noise model (Sect. 3.3.3). We fixed the inclination of planet c ($i_{\mathrm{c}}$) to 89.6 deg (Table 5), and assumed the external planet d to have an inclination of 90 deg. We also fixed the longitude of nodes ($\Omega$) of both planets to 0 deg. We then varied the ratio of the planetary to stellar mass ($M_{\rm p}/M_{\star}$), the period ($P$), the eccentricity ($e$) and the argument of periastron ($\omega$) as $\sqrt{e}\cos\omega$ and $\sqrt{e}\sin\omega$, and the mean longitude ($\lambda$666 the mean longitude is defined as $\lambda=\mathcal{M}+\omega+\Omega$, where $\mathcal{M}$ is the mean anomaly.), of both planets c and d. All the orbital parameters are osculating astrocentric parameters at the reference time $\mathrm{BJD_{TDB}}-2450000.0=7081.0$ and have uniform priors with physical boundaries, which are: $M_{\mathrm{c,d}}=\mathcal{U}(0.1,20)\ M_{\oplus}$, $P_{\mathrm{c}}=\mathcal{U}(42,48)$ d, $P_{\mathrm{d}}=\mathcal{U}(50,250)$ d, $e_{\mathrm{c,d}}=\mathcal{U}(0.0,0.5)$, $\omega_{\mathrm{c,d}}=\mathcal{U}(0.0,360)$ deg, $\lambda_{\mathrm{c,d}}=\mathcal{U}(0.0,360)$ deg. Firstly, we searched for an optimal configuration with a differential evolution algorithm (DE, Storn & Price 1997) implemented as PyDE within pytransit [^7] https://github.com/hpparvi/PyTransit (Parviainen 2015) with a population of 76 configurations that evolved for 76000 generations. Then, we passed the optimal configuration to the python code emcee (Foreman-Mackey et al. 2013, 2019). We used as walkers the same number of the DE population (76), and we mixed the sampler algorithm as in Nascimbeni et al. (2023). We ran emcee for 2.5 million steps, removed as burn-in the first 2 million, and used a thinning factor of 100. We computed as parameter uncertainties the high density interval (HDI) at $68.27\%$ [^8] This is the equivalent of the confidence intervals at the $16^{\mathrm{th}}$ and $84^{\mathrm{th}}$ percentile. of the posterior distribution, and our best-fit solution as the maximum likelihood estimator (MLE), that is, the configuration of parameters that maximizes the log-likelihood ($\log\mathcal{L}$) of the posterior distribution within the HDI. Table 9 gives the results of the best-fit parameters, and Figs. 8 and 9 show the best-fit models of TTVs and RVs, respectively. We checked the best-fit model assuming that planet d does not transit, and found an average difference on the transit times of less than 1 second. This is well below our sensitivity threshold given by the mean transit uncertainty of two minutes. Therefore, the effect of using an orbital inclination slightly different from 90 deg for planet d is negligible in our analysis.
@@
The parameters are in agreement with the different solutions of Sect.3. In particular, the eccentricities of both planets c and d are more precise than those retrieved in the DE-MCMC and MultiNest analyses (Sect. 3.3.1 and 3.3.2). We decided to test a model with the same number of bodies, but fixing $e_{\mathrm{d}}=0$, and we ran PyDE and emcee with the same number of walkers (76), but with a lower number of steps (76000 PyDE generations, 1M emcee steps with 400000 as burn-in, and a further thinning factor of 100). We computed the BIC for the eccentric (ecc) and for the circular (circ) best-fit model, and we found that the eccentric model is strongly favored: $\Delta\mathrm{BIC}=\mathrm{BIC(ecc)}-\mathrm{BIC(circ)}=-13$. (e.g., Kass & Raftery 1995).
@@
Table 4: Summary of the best-fit MLE parameters for the Kepler-10 (c + d) system analysed with TRADES.
@@
 [^9] The columns give the name of the parameter, its best-fit value with its associated $68.27\%$ HDI. Astrocentric orbital parameters are defined at $2450000.07081.0\,\mathrm{BJD_{TDB}}-$ 2450000.0. $\mathcal{M}$ is the mean anomaly computed as $\mathcal{M}=\lambda-\omega-\Omega$. We fixed the value of $i_{\mathrm{d}}=90$ deg, and so $M_{\mathrm{p}}$ of planet d is the minimum mass. See Section 3.5 for details.
Figure 8: Kepler-10 c Observed-Calculated (O-C) diagram, defined as in Fig. 3. Upper panel: observed (orange circles) TTVs, the best-fit model (black circles and line), and the gray areas are the 1, 2, and $3\sigma$ uncertainty (from darker to lighter) of 100 samples drawn from the posterior distribution. Lower panel: residuals between the observed $T_{c}$ and simulated ones with TRADES.
Figure 9: Kepler-10 RVs from the TRADES analysis after subtracting the signal of Kepler-10 b. Upper panel: observed RVs (removed RV offset for each data-set) as colored circles (magenta for HN-1 and blue for HN-2), and the TRADES best-fit model (black line). Lower panel: residuals between the observed RVs and simulated ones with TRADES.
@@ 4 Planet parameters
We updated the orbital and physical parameters of the Kepler-10 planets in Table 5, based on i) the stellar parameters in D14; ii) the results of the DE-MCMC transit and RV modeling (Sect. 3), also for consistency with both D14 and Bonomo et al. (2023) (we recall that fully compatible results were obtained with the other two techniques employed; see Table LABEL:table_harpsn_param); and iii) the combined analysis of TTVs and RVs with TRADES for planets c and d.
@@
For the outer non-transiting planet Kepler-10 d we found a more accurate orbital period of $P=151.06\pm 0.48$ d from the HARPS-N RVs (Table 5), which is consistent with both the analysis of TTVs (Sect. 3.1.1) and the combined analysis of TTVs and RVs (cf. Sect. 3.5 and Table 5), and derived a minimum mass of $M_{\rm p,d}\sin{i}=12.00\pm 2.15$ $\mathrm{M}_{\oplus}$
@@
Orbital semi-major axis $a$ [AU] $0.01685\pm 0.00013$ D14
@@
Orbital semi-major axis $a$ [AU] $0.2410\pm 0.0019$ D14
@@
Orbital semi-major axis $a$ [AU] $0.5379\pm 0.0043$ This work
@@ 5 Discussion and conclusions
The innermost planet Kepler-10 b is a rocky super-Earth that likely formed inside the water snowline at $\sim 1-3$ AU, i.e. in a possibly dry environment, and then migrated inward to the current semi-major axis where it is observed now. The planet may possibly have lost a primordial atmospheric H/He envelope because of the strong XUV stellar irradiation received in the first $\sim 100$ Myr (e.g., Owen & Wu 2013), but may have retained a thin, recently formed collisional secondary atmosphere (Singh et al. 2022). Since Kepler-10 b is located on the pure-silicate isocomposition curve (Fig. 10), its core mass fraction is broadly consistent with zero, though a small iron core may be present.
On the contrary, Kepler-10 c, by its position in Fig. 10 and its low equilibrium temperature ($T_{\rm eq}\simeq 580$ K), may be a water world, with a significant ($\sim 40-70\%$) mass fraction of H2O in a differentiated or miscible interior (e.g., Luo et al. 2024). Unlike many of the exoplanets in the figure, which lie above the pure-H2O (“Ices”) mass-radius curve due to the presence of some atmospheric envelope, Kepler-10 c lies below it, despite being on the right-hand-side of the Cosmic Hydrogen and Ice Loss Line, where atmospheric loss due to stellar radiation is expected to be negligible (cf. Zeng & Jacobsen 2024). The existence of water worlds is predicted by models of planet formation and migration (e.g., Venturini et al. 2024; Burn et al. 2024). Kepler-10 c may thus join the fifteen or so possible water worlds orbiting relatively far from their FGK dwarf hosts, and hence having rather low equilibrium temperatures $T_{\rm eq}\leq 600$ K, located in between the $100\%$ silicate and $100\%$ water isocomposition curves in Fig. 11 (see also Luque & Pallé 2022 for possible water worlds around M dwarfs).
Figure 10: Mass-radius diagram of small ($R_{\rm p}\leq 4\leavevmode\nobreak\ \hbox{$\mathrm{R}_{\oplus}$}\,$) planets with mass and radius determinations better than $4\sigma$ and $10\sigma$, respectively, color-coded by planet equilibrium temperatures. The different solid curves, from bottom to top, correspond to planet compositions of $100\%$ iron, $33\%$ iron core and $67\%$ silicate mantle (Earth-like composition), $100\%$ silicates, $50\%$ rocky interior and $50\%$ water, $100\%$ water, rocky interiors and $1\%$ or $2\%$ hydrogen-dominated atmospheres (Zeng & Sasselov 2013). The gray dark circles indicate Venus (V), the Earth (E), Uranus (U), and Neptune (N). Kepler-10 b and c are indicated with squares.
Figure 11: Same as Fig. 10 for planets orbiting FGK dwarfs, with color-coded equilibrium temperatures $T_{\rm eq}\leq 600$ K. Kepler-10 c is indicated with a square.
@@
The rather large density of Kepler-10 c is in agreement with the trend that sub-Neptunes that are relatively far from MMR101010according to the criterion given in Leleu et al. (2024). are denser (Leleu et al. 2024). This may provide an additional potential scenario to explain its density: the ejection of all or most of its atmosphere in post-disc dynamical instabilities (e.g., Izidoro et al. 2017).
Kepler-10 d might be similar to Kepler-10 c in terms of its composition, its minimum mass being consistent with the mass of Kepler-10 c within $1\sigma$. Kepler-10 d might thus potentially be a water world or have a composition more similar to Uranus and Neptune, i.e. with a larger H/He envelope, given that it is colder and likely formed farther away. In any case, the lack of the measurement of both its radius and true mass makes any discussion quite speculative.
Given the long RV time series of Kepler-10 spanning $\sim 11$ years, we estimated the sensitivity of our data (completeness) to the presence of possible outer cold Jupiters as in Bonomo et al. (2023) (see their Sect. 3.3). The completeness is shown in Fig. 12, where yellow cells indicate 100% detectability of the cold Jupiter, lighter to green cells show decreasing detection probability, and the blue ones correspond to 0% detectability. It is clear that there are no giant planets orbiting the Kepler-10 star within 10 AU, as for the great majority of Kepler and K2 small planet systems followed by HARPS-N (Bonomo et al. 2023); otherwise, we would have easily detected it. This is consistent with the scenario that no giant planet could act as a dynamical barrier to the inward migration of the sub-Neptunes Kepler-10 c and Kepler-10 d (e.g., Izidoro et al. 2015).
Figure 12: Completeness map of Kepler-10, showing the sensitivity of the HARPS-N RV measurements to the presence of possible cold Jupiters through experiments of injection and retrieval of Doppler signals. The detectability inside the cell is indicated by its color, according to the color-bar on the right side of the figure.
This work highlights the crucial importance of the intensive long-term RV monitoring of planetary systems with relatively long-period ($P\gtrsim 30$ d) transiting small planets, also in view of the forthcoming PLATO space mission (Rauer et al. 2014, 2024). As in the case of Kepler-10, even several hundreds of RV measurements may be needed to i) precisely determine the masses and hence bulk densities of the transiting planets and ii) properly characterize the system architecture by searching for additional non-transiting small and giant planets in outer regions. The increasing number of planetary systems better characterized in this way will lead to a more comprehensive understanding of the formation and evolution of sub-Neptunes and super-Earths.
@@ Appendix A Generalized Lomb-Scargle Periodograms of the HARPS-N radial velocities
Figure 13: GLS periodograms of the 291 HARPS-N radial velocities and residuals as a function of frequency in cycles/day (or day-1). From top to bottom are displayed the periodograms of radial velocities; residuals after subtracting the signal of Kepler-10 b; residuals after subtracting the signals of both Kepler-10 b and c; residuals after subtracting the signals of Kepler-10 b, c, and d; and the spectral window. The dotted vertical lines indicate the orbital periods of planets b (green), c (red), and d (light blue). The horizontal dash-dotted lines show the False Alarm Probability level of $10^{-5}$.
@@ B.1 False inclusion probability
This appendix gives the details of the analysis of the HARPS-N data on Kepler-10 processed with YARARA (Cretignier et al. 2021). We define the true inclusion probability (TIP) as the posterior probability of the following event: For a given range of periods $I$, there is at least one planet with period $P\in I$. By analogy with the FAP, we also define the false inclusion probability (FIP), which is the probability that there is no planet with period in interval $I$. Formally, denoting the data by $y$, the TIP and FIP are
@@
We perform the FIP calculation up to four planets with two different types of priors for the semi-amplitude. The rationale is that, as we have shown in Hara et al. (2022b), the significance of low-amplitude signals highly depends on the on the chosen prior. In the two sets of priors, we use a $\log$-uniform prior on period, we constrain the periods and phase of two of the planets according to the constraints from the transit. The eccentricity has a Beta prior as in Kipping (2014), the prior on the argument of periastron is uniform.
@@
We then compute the FIP/TIP as the probability to have at least one planet with frequency in the interval $[f-1/T_{\mathrm{obs}},f+1/T_{\mathrm{obs}}]$ where $T_{\mathrm{obs}}$ is the total time of observation, for an array of equispaced frequencies. In the first set of priors, we use a log-uniform prior from 0.1 to 20  m s-1on the RV semi-amplitudes, and obtain the FIP periodogram shown in Fig. 5.
@@
In the second set of priors, the prior is not defined on semi-amplitude but on the mass. This means that for each set of period, eccentricity and time of passage at periastron (or equivalently, mean anomaly at a reference time), the prior on semi-amplitude is scaled to correspond to the same prior on mass. This choice implicitly supposes that the probability of occurrence of a given type of planet (Super-Earth, Neptune) does not depend on semi-major axis. Since, for a given mass, the semi-amplitude decreases with separation, this will also have the effect to boost the significance of low amplitude signals. We use a Gaussian mixture model for the prior on mass, which follows a Rayleigh prior with $\sigma=5M_{\oplus}$ or $\sigma=15M_{\oplus}$ with a probability 1/2. With this choice, we can marginalise the likelihood analytically over the linear parameters as shown in Hara et al. (2022b), Appendix C. In that case, a 68 days signal reaches a FIP of 4% (See Fig. 5, b). Note that the two different priors are applied to two slightly different versions of the YARARA processing of RVs.
@@
We did the same analyses as before but including a red noise term with an exponential kernel (free jitter and time-scale). The corresponding FIP periodograms are shown in Fig. 14, where the 150.7d signal is much less significant and completely disappears, respectively. However, we caution that red noise model is expected to damp the amplitude of long period signals, as shown in Delisle et al. (2020). It is common for this type of noise model to suppress viable planets. If what happens at low frequency was truly red noise, then we would expect the white noise model to find several unclear planet candidates at long periods, due to a diffuse power at low frequencies. However in Kepler-10, the white noise model shows a clear candidate with a well defined period. As a consequence, the interpretation of the FIP periodogram with a red noise model with exponential kernel is subject to caution.
@@
Figure 14: FIP periodograms of the Kepler-10 HARPS-N data processed with YARARA for a red noise model, for different priors on semi amplitude. (a) is obtained with a log-uniform prior on semi-amplitude, (b) with a Gaussian mixture model on mass. In yellow we represent the true inclusion probability (TIP), the probability of having a planet in the range $[\omega-\Delta\omega,\omega+\Delta\omega]$ as a function of $\omega$. In blue, we represent $-\log_{10}FIP$, where FIP = 1 - TIP is the false inclusion probability.
In Table 7 and Table 7. We show the evidences $\mathcal{Z}$ and uncertainty on them (standard deviation of three runs) for the log-uniform prior on K, white and red noise assumptions, respectively. We can now compute the TIP of the 150.7 days planet averaged over the noise model. Denoting by $I$ the frequency interval $[\frac{2\pi}{150.7}-\Delta\omega,\frac{2\pi}{150.7}+\Delta\omega]$
$p(\omega\in I\mid y)=p(\omega\in I\mid W,y)p(W\mid y)+p(\omega\in I\mid R,y)p(R\mid y).$ (6)
where $W$ and $R$ represent the white and red noise signal. The two terms $p(\omega\in I\mid W,y)$ and $p(\omega\in I\mid R,y)$ are equal to 1-FIP of the 150.7 days signal for the white and red noise models respectively. We have $p(\omega\in I\mid R,y)=0$. Assuming white and red noise have both a prior probability of 1/2, we have
$p(W\mid y)=\frac{p(y\mid W)}{p(y\mid W)+p(y\mid R)}$ (7)
and
$p(y\mid W)=\sum_{k=0}^{3}p(y\mid W,k)p(k)$ (8)
$p(y\mid W)$ $=\operatorname{e}^{-732.70}/4$ (9)
Which means that $p(W\mid y)=0.84)$. As a result, the probability to have a signal at 150.7 days is $(1-10^{-2.16})\times 0.84\sim 0.84$. The probability that there is a signal in $I$ marginalized on the noise models considered here is 84%. Conversely, we have a FIP for the 150.7 days signal of 1 - 0.84 = 0.16.

### v2

@@ 3.6 Simultaneous modeling of radial velocities and transit timing variations
The TRADES code cannot manage complex stellar activity, it can only fit an RV offset ($\gamma$) and a jitter term ($\log_{2}{\sigma_{\rm j}}$) for each RV dataset; this is not an issue in our case, given the low level of stellar activity and the fact that the planet parameters do not critically depend on the noise model (Sect. 3.4.3). We fixed the inclination of planet c ($i_{\mathrm{c}}$) to 89.6 deg (Table 5), and assumed the external planet d to have an inclination of 90 deg. We also fixed the longitude of nodes ($\Omega$) of both planets to 0 deg. We then varied the ratio of the planetary to stellar mass ($M_{\rm p}/M_{\star}$), the period ($P$), the eccentricity ($e$) and the argument of periastron ($\omega$) as $\sqrt{e}\cos\omega$ and $\sqrt{e}\sin\omega$, and the mean longitude ($\lambda$666 the mean longitude is defined as $\lambda=\mathcal{M}+\omega+\Omega$, where $\mathcal{M}$ is the mean anomaly.), of both planets c and d. All the orbital parameters are osculating astrocentric parameters at the reference time $\mathrm{BJD_{TDB}}-2450000.0=7081.0$ and have uniform priors with physical boundaries, which are: $M_{\mathrm{c,d}}=\mathcal{U}(0.1,20)\ M_{\oplus}$, $P_{\mathrm{c}}=\mathcal{U}(42,48)$ d, $P_{\mathrm{d}}=\mathcal{U}(50,250)$ d, $e_{\mathrm{c,d}}=\mathcal{U}(0.0,0.5)$, $\omega_{\mathrm{c,d}}=\mathcal{U}(0.0,360)$ deg, $\lambda_{\mathrm{c,d}}=\mathcal{U}(0.0,360)$ deg. Firstly, we searched for an optimal configuration with a differential evolution algorithm (DE, Storn & Price 1997) implemented as PyDE within pytransit [^7] https://github.com/hpparvi/PyTransit (Parviainen 2015) with a population of 76 configurations that evolved for 76000 generations. Then, we passed the optimal configuration to the python code emcee (Foreman-Mackey et al. 2013, 2019). We used as walkers the same number of the DE population (76), and we mixed the sampler algorithm as in Nascimbeni et al. (2023). We ran emcee for 2.5 million steps, removed as burn-in the first two million, and used a thinning factor of 100. We computed as parameter uncertainties the high density interval (HDI) at $68.27\%$ [^8] This is the equivalent of the confidence intervals at the $16^{\mathrm{th}}$ and $84^{\mathrm{th}}$ percentile. of the posterior distribution, and our best-fit solution as the maximum likelihood estimator (MLE), that is, the configuration of parameters that maximizes the log-likelihood ($\log\mathcal{L}$) of the posterior distribution within the HDI. Table 9 gives the results of the best-fit parameters, and Figs. 8 and 9 show the best-fit models of TTVs and RVs, respectively. We checked the best-fit model assuming that planet d does not transit, and found an average difference on the transit times of less than one second. This is well below our sensitivity threshold given by the mean transit uncertainty of two minutes. Therefore, the effect of using an orbital inclination slightly different from 90 deg for planet d is negligible in our analysis.
@@
The parameters are in agreement with the different solutions of Sect.3. In particular, the eccentricities of both planets c and d are more precise than those retrieved in the DE-MCMC and MultiNest analyses (Sects. 3.4.1 and 3.4.2). We decided to test a model with the same number of bodies, but fixing $e_{\mathrm{d}}=0$, and we ran PyDE and emcee with the same number of walkers (76), but with a lower number of steps (76000 PyDE generations, one million emcee steps with 400000 as burn-in, and a further thinning factor of 100). We computed the BIC for the eccentric (ecc) and for the circular (circ) best-fit model, and we found that the eccentric model is strongly favored: $\Delta\mathrm{BIC}=\mathrm{BIC(ecc)}-\mathrm{BIC(circ)}=-13$. (e.g., Kass & Raftery 1995).
@@
Table 4: Best-fit MLE parameters for the Kepler-10 (c + d) system analyzed with TRADES.
@@
 [^9] The columns give the name of the parameter, its best-fit value with its associated $68.27\%$ HDI. Astrocentric orbital parameters are defined at $2450000.07081.0\,\mathrm{BJD_{TDB}}-$ 2450000.0. The mean anomaly $\mathcal{M}$ is computed as $\mathcal{M}=\lambda-\omega-\Omega$. We fixed the value of $i_{\mathrm{d}}=90$ deg, and so the $M_{\mathrm{p}}$ of planet d is the minimum mass. See Sect. 3.6 for details.
Figure 8: Kepler-10 c O-C diagram, defined as in Fig. 3. Top panel: Observed (orange circles) TTVs, the best-fit model (black circles and line), and the gray areas are the 1, 2, and $3\sigma$ uncertainty (from darker to lighter) of 100 samples drawn from the posterior distribution. Bottom panel: Residuals between the observed $T_{c}$ and simulated ones with TRADES.
Figure 9: Kepler-10 RVs from the TRADES analysis after subtracting the signal of Kepler-10 b. Top panel: Observed RVs (removed RV offset for each dataset) as colored circles (magenta for HN-1 and blue for HN-2), and the TRADES best-fit model (black line). Bottom panel: Residuals between the observed RVs and simulated ones with TRADES.
@@ 4 Planet parameters
We updated the orbital and physical parameters of the Kepler-10 planets in Table 5, based on (i) the stellar parameters in D14; (ii) the results of the DE-MCMC transit and RV modeling (Sect. 3), also for consistency with both D14 and Bonomo et al. (2023) (we recall that fully compatible results were obtained with the other two techniques employed; see Table LABEL:table_harpsn_param); and (iii) the combined analysis of TTVs and RVs with TRADES for planets c and d.
@@
For the outer non-transiting planet Kepler-10 d, we found a more accurate orbital period of $P=151.06\pm 0.48$ d from the HARPS-N RVs (Table 5), which is consistent with both the analysis of TTVs (Sect. 3.2) and the combined analysis of TTVs and RVs (cf. Sect. 3.6 and Table 5), and derived a minimum mass of $M_{\rm p,d}\sin{i}=12.00\pm 2.15$ $\mathrm{M}_{\oplus}$
@@
Orbital semimajor axis $a$ [AU] $0.01685\pm 0.00013$ D14
@@
Orbital semimajor axis $a$ [AU] $0.2410\pm 0.0019$ D14
@@
Orbital semimajor axis $a$ [AU] $0.5379\pm 0.0043$ This work
@@ 5 Discussion and conclusions
Figure 10: Left panel: Mass-radius diagram of small ($R_{\rm p}\leq 4\leavevmode\nobreak\ \hbox{$\mathrm{R}_{\oplus}$}\,$) planets with mass and radius determinations better than $4\sigma$ and $10\sigma$, respectively, color-coded by planet equilibrium temperatures. The different solid curves, from bottom to top, correspond to planet compositions of $100\%$ iron, $33\%$ iron core and $67\%$ silicate mantle (Earth-like composition), $100\%$ silicates, $50\%$ rocky interior and $50\%$ water, $100\%$ water, rocky interiors and $1\%$ or $2\%$ hydrogen-dominated atmospheres (Zeng & Sasselov 2013). The gray dark circles indicate Venus (V), the Earth (E), Uranus (U), and Neptune (N). Right panel: Same as left panel for planets orbiting FGK dwarfs, with color-coded equilibrium temperatures $T_{\rm eq}\leq 600$ K. Kepler-10 b and c are indicated with squares.
The innermost planet Kepler-10 b is a rocky super-Earth that likely formed inside the water snowline at $\sim 1-3$ AU, that is, in a possibly dry environment, and then migrated inward to the current semimajor axis where it is observed now. The planet may possibly have lost a primordial atmospheric H/He envelope because of the strong XUV stellar irradiation received in the first $\sim 100$ Myr (e.g., Owen & Wu 2013), but may have retained a thin, recently formed collisional secondary atmosphere (Singh et al. 2022). Since Kepler-10 b is located on the pure-silicate isocomposition curve (Fig. 10, left panel), its core mass fraction is broadly consistent with zero, though a small iron core may be present.
On the contrary, Kepler-10 c, by its position in Fig. 10 and its low equilibrium temperature ($T_{\rm eq}\simeq 580$ K), may be a water world, with a significant ($\sim 40-70\%$) mass fraction of H2O in a differentiated or miscible interior (e.g., Luo et al. 2024). Unlike many of the exoplanets in the figure, which are above the pure-H2O mass-radius curve due to the presence of some atmospheric envelope, Kepler-10 c lies below it, despite being on the right-hand side of the Cosmic Hydrogen and Ice Loss Line, where atmospheric loss due to stellar radiation is expected to be negligible (cf. Zeng & Jacobsen 2024). The existence of water worlds is predicted by models of planet formation and migration (e.g., Venturini et al. 2024; Burn et al. 2024). Kepler-10 c may thus join the fifteen or so possible water worlds that orbit relatively far from their FGK dwarf hosts, and hence have rather low equilibrium temperatures $T_{\rm eq}\leq 600$ K, located in between the $100\%$ silicate and $100\%$ water isocomposition curves (Fig. 10, right panel; see also Luque & Pallé 2022 for possible water worlds around M dwarfs).
@@
The rather large density of Kepler-10 c is in agreement with the trend that sub-Neptunes that are relatively far from MMR101010according to the criterion given in Leleu et al. (2024). are denser (Leleu et al. 2024). This may provide an additional potential scenario to explain its density: the ejection of all or most of its atmosphere in post-disk dynamical instabilities (e.g., Izidoro et al. 2017).
Kepler-10 d might be similar to Kepler-10 c in terms of its composition, its minimum mass being consistent with the mass of Kepler-10 c within $1\sigma$. Kepler-10 d might thus potentially be a water world or have a composition more similar to Uranus and Neptune, that is, with a larger H/He envelope, given that it is colder and likely formed farther away. In any case, the lack of the measurement of both its radius and true mass makes any discussion quite speculative.
Given the long RV time series of Kepler-10 spanning $\sim 11$ years, we estimated the sensitivity of our data (completeness) to the presence of possible outer cold Jupiters as in Bonomo et al. (2023, see their Sect. 3.3). The completeness is shown in Fig. 11, where yellow cells indicate 100% detectability of the cold Jupiter, lighter to green cells show decreasing detection probability, and the blue ones correspond to 0% detectability. It is clear that there are no giant planets orbiting the Kepler-10 star within 10 AU, as is the case for the great majority of Kepler and K2 small planet systems followed by HARPS-N (Bonomo et al. 2023); otherwise, we would have easily detected it. This is consistent with the scenario that no giant planet could act as a dynamical barrier to the inward migration of the sub-Neptunes Kepler-10 c and Kepler-10 d (e.g., Izidoro et al. 2015).
Figure 11: Completeness map of Kepler-10, showing the sensitivity of the HARPS-N RV measurements to the presence of possible cold Jupiters through experiments of injection and recovery of Doppler signals. The detectability inside the cell is indicated by its color, according to the colorbar on the right-hand side of the figure.
This work highlights the crucial importance of intensive long-term RV monitoring of planetary systems with relatively long-period ($P\gtrsim 30$ d) transiting small planets, also in view of the forthcoming PLAnetary Transits and Oscillations of stars (PLATO) space mission (Rauer et al. 2014, 2024). As in the case of Kepler-10, even several hundreds of RV measurements may be needed to (i) precisely determine the masses and hence bulk densities of the transiting planets, and (ii) properly characterize the system architecture by searching for additional non-transiting small and giant planets in outer regions. The increasing number of planetary systems better characterized in this way will lead to a more comprehensive understanding of the formation and evolution of sub-Neptunes and super-Earths.
@@ Appendix A Generalized Lomb-Scargle periodograms of the HARPS-N radial velocities
Figure 12: Generalized Lomb-Scargle periodograms of the 291 HARPS-N radial velocities and residuals as a function of frequency in cycles/day (or day-1). From top to bottom are displayed the periodograms of radial velocities; residuals after subtracting the signal of Kepler-10 b; residuals after subtracting the signals of both Kepler-10 b and c; residuals after subtracting the signals of Kepler-10 b, c, and d; and the spectral window. The dotted vertical lines indicate the orbital periods of planets b (green), c (red), and d (light blue). The horizontal dash-dotted lines show the false alarm probability level of $10^{-5}$.
@@ B.1 False inclusion probability
This appendix gives the details of the analysis of the HARPS-N data on Kepler-10 processed with YARARA-v2 (Cretignier et al. 2021, 2022). We define the true inclusion probability (TIP) as the posterior probability of the following event: for a given range of periods $I$, there is at least one planet with period $P\in I$. By analogy with the FAP, we also define the false inclusion probability (FIP), which is the probability that there is no planet with period $P$ in interval $I$. Formally, denoting the data by $y$, the TIP and FIP are
@@
We perform the FIP calculation up to four planets with two different types of priors for the semiamplitude. The rationale is that, as we have shown in Hara et al. (2022b), the significance of low-amplitude signals highly depends on the chosen prior. In the two sets of priors, we use a $\log$-uniform prior on period, we constrain the periods and phase of two of the planets according to the constraints from the transit. The eccentricity has a Beta prior as in Kipping (2014), the prior on the argument of periastron is uniform.
@@
We then compute the FIP/TIP as the probability to have at least one planet with frequency in the interval $[f-1/T_{\mathrm{obs}},f+1/T_{\mathrm{obs}}]$ where $T_{\mathrm{obs}}$ is the total time of observation, for an array of equispaced frequencies. In the first set of priors, we use a log-uniform prior from 0.1 to 20  m s-1on the RV semiamplitudes, and obtain the FIP periodogram shown in Fig. 5 (a).
@@
In the second set of priors, the prior is not defined on semiamplitude but on the mass. This means that, for each set of period, eccentricity and time of passage at periastron (or equivalently, mean anomaly at a reference time), the prior on semiamplitude is scaled to correspond to the same prior on mass. This choice implicitly supposes that the probability of occurrence of a given type of planet (super-Earth, Neptune) does not depend on semimajor axis. Since, for a given mass, the semiamplitude decreases with separation, this will also have the effect to boost the significance of low amplitude signals. We use a Gaussian mixture model for the prior on mass, which follows a Rayleigh prior with $\sigma=5M_{\oplus}$ or $\sigma=15M_{\oplus}$ with a probability 1/2. With this choice, we can marginalize the likelihood analytically over the linear parameters as shown in Appendix C of Hara et al. (2022b). In that case, a 68 d signal reaches a FIP of 4% (see Fig. 5, b).
@@
We did the same analyses as before but including a red noise term with an exponential kernel (free jitter and time-scale). The corresponding FIP periodograms are shown in Fig. 13, where the 150.7 d signal is much less significant and completely disappears, respectively. However, we caution that the red noise model is expected to damp the amplitude of long-period signals, as shown in Delisle et al. (2020). It is common for this type of noise model to suppress viable planets. If what happens at low frequency was truly red noise, then we would expect the white noise model to find several unclear planet candidates at long periods, due to a diffuse power at low frequencies. However, in Kepler-10 the white noise model shows a clear candidate with a well defined period. As a consequence, the interpretation of the FIP periodogram with a red noise model with exponential kernel is subject to caution.
@@
Figure 13: False inclusion probability periodograms of the Kepler-10 HARPS-N data processed with YARARA-v2 for a red noise model, for different priors on semiamplitude. (a) is obtained with a log-uniform prior on semiamplitude, (b) with a Gaussian mixture model on mass. In yellow we represent the true inclusion probability (TIP), the probability of having a planet in the range $[\omega-\Delta\omega,\omega+\Delta\omega]$ as a function of $\omega$. In blue, we represent $-\log_{10}FIP$, where FIP = 1 - TIP is the false inclusion probability.
@@
In Table 7 and Table 7 we show the evidences $\mathcal{Z}$ and uncertainty on them (standard deviation of three runs) for the log-uniform prior on K, white and red noise assumptions, respectively. We can now compute the TIP of the 150.7 d planet averaged over the noise model. Denoting by $I$ the frequency interval $[\frac{2\pi}{150.7}-\Delta\omega,\frac{2\pi}{150.7}+\Delta\omega]$, we have
$p(\omega\in I\mid y)=p(\omega\in I\mid W,y)p(W\mid y)+p(\omega\in I\mid R,y)p(R\mid y),$ (6)
where $W$ and $R$ represent the white and red noise signal. The two terms $p(\omega\in I\mid W,y)$ and $p(\omega\in I\mid R,y)$ are equal to 1-FIP of the 150.7 d signal for the white and red noise models, respectively. We have $p(\omega\in I\mid R,y)=0$. Assuming white and red noise have both a prior probability of 1/2, we have
$p(W\mid y)=\frac{p(y\mid W)}{p(y\mid W)+p(y\mid R)}$ (7)
and
$p(y\mid W)=\sum_{k=0}^{3}p(y\mid W,k)p(k)$ (8)
$p(y\mid W)$ $=\operatorname{e}^{-732.70}/4$ (9)
which means that $p(W\mid y)=0.84$. As a result, the probability to have a signal at 150.7 d is $(1-10^{-2.16})\times 0.84\sim 0.84$. The probability that there is a signal in $I$ marginalized on the noise models considered here is 84%. Conversely, we have a FIP for the 150.7 d signal of 1 - 0.84 = 0.16.